OpenStudy (anonymous):

If the graph of f(x) = x2 is shifted up 9 units, then what would be the equation of the new graph?

OpenStudy (anonymous):

\[f(x)=x^2\] ?

OpenStudy (shayaan_mustafa):

\[f(x)=x ^{2}+9\]

OpenStudy (anonymous):

Reason like this, what happens at x=0? f(0)=0^2

OpenStudy (shayaan_mustafa):

@Noliec you mean if x=0 then it could be x=9?

OpenStudy (anonymous):

You want the curve to be 9 at f(0) because you're moving it up 9 units. This leaves the equation f(0)=9 --> f(x)=x^2+9

OpenStudy (shayaan_mustafa):

yes that's right.

OpenStudy (shayaan_mustafa):

Noliec is right.